OpenStudy (anonymous):

Help me solve: sin^2x-sinx=0

OpenStudy (dan815):

0 and 1 sin x=0 and 1

OpenStudy (dan815):

npi, pi/2+n2pi

OpenStudy (swissgirl):

\[\sin^2x-\sin x=0\] \[\sin x( \sin x-1)=0\] \[\frac{\sin x( \sin x-1)}{\sin x}=\frac{0}{\sin x}\] \[ \sin x-1=0\] \[ \sin x=1\] \[\sin^{-1} (1)=x\] This occurs at x=90

OpenStudy (swissgirl):

\[\sin x=0\] \[\sin ^{-1}(0)=x\] This occurs at x=0 and x=360

OpenStudy (swissgirl):

dang it ill give ya a medal anyways ARGHHHHHH I HATE YOUUUUUUUUUUUUUUU

OpenStudy (dan815):

sin^2x-sinx=0 let y= sin x y^2-y=0 y(y-1)=0, y=0 or 1 y=sinx 0=sinx, x=n*pi 1=sinx, x=pi/2+2pi*n

OpenStudy (dan815):

dont forget osscilations

OpenStudy (dan815):

90+n*360. nEall integers

OpenStudy (swissgirl):

dan what does n stand for?

OpenStudy (swissgirl):

ohhh for alll cycles

OpenStudy (swissgirl):

gotcha gotcha

OpenStudy (dan815):

and for every 180 degree angle so 180 + n 180

OpenStudy (dan815):

or npi

OpenStudy (dan815):

lose the habit of using degrees

OpenStudy (swissgirl):

hahaha ok

OpenStudy (swissgirl):

fuk u

OpenStudy (dan815):

please dont swear on OS

OpenStudy (dan815):

It is a user friendly site for all ages

OpenStudy (swissgirl):

hahahahaha

OpenStudy (dan815):

welcome to OS sloan

OpenStudy (swissgirl):

Im gonna beat u up soon

OpenStudy (swissgirl):

@sloan16 fell asleep -.-

OpenStudy (anonymous):

thank you guys so much

OpenStudy (dan815):

You are welcome sloan

OpenStudy (dan815):

give a medal to swissgirl

OpenStudy (swissgirl):

dan shut up -.-